OpenStudy (anonymous):

15. Suppose the polynomial function below represents the power generated by a wind turbine, where x represents the wind speed in meters per second and y represents the kilowatts generated. Interpret ƒ(10).

OpenStudy (anonymous):

what is the function given?

OpenStudy (anonymous):

ƒ(x) = 0.08x^3 + x^2 + x + 0.26

OpenStudy (anonymous):

use you calculator to find f(10) ƒ(10) = 0.08(10)^3 + (10)^2 + 10 + 0.26

OpenStudy (anonymous):

ƒ(x) = 0.08x^3 + x^2 + x + 0.26 = (0.08 *10^3) + 10^2 + 10 + 0.26 =80+10^2+10+.26 =80+100+10+.26 =190.26 Kilowatts